# Ensembled Direct Multi Step forecasting methodology with comparison on macroeconomic and financial data

### v1

@@ 3 Ensembled Direct Multi-Step forecasting
In this section, we will describe the Ensembled Direct Multi-Step (EDMS) forecasting method in detail. Subsection 2.1 presents the models used for the ensemble forecasts, Subsection 2.2 discusses the ensembling method itself, and finally, Subsection 2.3 explores the application of Direct Multi-Step (DMS) forecasting combined with the ensembling method.
@@ 3.1 Models
To enhance the modelling capability of the ensemble, we carefully selected a diverse set of models. We considered only those models that perform well in terms of accuracy and computational efficiency. Additionally, we included models across the full spectrum of complexity, ranging from simple to highly complex. For the simplest models, we employed basic forecasting using the average of past growth rates, as well as Linear and Polynomial Regression. For models of medium complexity, we utilized classical time series models, specifically General Exponential Smoothing (ETS). For the most complex models, we implemented an advanced deep learning model, namely LSTM. In selecting these models, we aimed to minimize the computing resources and time required for forecasting, as economic and financial datasets tend to be relatively large. At the same time, we strive to maintain the highest possible accuracy. Prophet was excluded from the set of models due to its significant computation time and minimal, if any, improvement in accuracy. We also have a constraint that the datasets must be sufficiently large, as a considerable amount of data is necessary to train more complex models such as LSTM.
@@
    Performance forecast - This involves forecasts generated using an 80/20 data split. These forecasts are calculated to assess the performance of the methods employed for computing weights in the ensembling process. We allocate 80% of the data for training and 20% for testing. Forecasts are then computed using the model trained on the training dataset. This split ratio was selected to ensure that there is sufficient data for effective model training, while also maintaining a reasonably large test set for proper model comparison. Experiments have demonstrated that the 80/20 ratio yields favourable results.
@@
    Ensembled forecast - Forecasts are generated using a 100/0 data split. This approach is used to compute the ensembled forecast. We utilize 100% of the data for training and perform forecasting with the model trained on the complete dataset in order to get the best possible forecast.
Let us now describe individually each model that is used in the ensemble:
@@
This is a type of Recurrent Neural Network (RNN) model. RNN is a natural choice for modelling time series due to its ability to effectively recognize and learn patterns in sequential data. Notably, LSTM is well-known and widely used model for this purpose, as outlined in [62]. We chose this model for several reasons. First, according to a literature review in [50], LSTM is the most extensively studied deep learning model for time series forecasting. Another reason for its selection is performance. Multiple sources indicate that LSTM outperforms widely used classical model ARIMA, as noted in [58, 44, 31, 72]. Furthermore, LSTM often prove to be the best performers among deep learning models, particularly for economic and financial data, as seen in [72, 59].
  The Long Short-Term Memory ANN architecture consists of sequentially connected modules with the number $n$ of these modules equal to the size of the input sequence $(y_{1},y_{2},\ldots,y_{n})$. In the sequence of modules, each module takes the following inputs:
@@
  • Previous cell state $C_{t-1}$ - the cell state of the previous module, which represents long term memory.
@@
The module is composed of the following components:
@@
    where $U_{f},W_{f}$ are weight matrices.
@@
    Update gate - this gate first computes a fully connected layer with hyperbolic tangent activation tanhtanh that proposes information for the new cell state $\widetilde{C}_{t}$, where
    $\tanh(y)=\frac{\exp(y)-\exp(-x)}{\exp(x)+\exp(-x)}.$
    Next, a fully connected layer with sigmoid activation $\sigma$, known as the input layer, selects which of the proposed information to add to the cell state. Both layers use the previous output $\widetilde{y}_{t}$ and the element of the input sequence $y_{t}$ as inputs. We then combine these two outputs to create an update to the state $C_{t}$. Those computations can be expressed as follows:
@@
    where $U_{i},W_{i}$ and $U_{C},W_{C}$ are weight matrices.
@@
    Output gate - this final gate consists of a fully connected layer with sigmoid activation $\sigma$ which selects which parts of the cell state to output and pass on:
@@
    where $U_{o},W_{o}$ are weight matrices.
The dimensions of $y_{t},\widetilde{y}_{t}$ are determined by the specific case at hand, which also determines the dimensions of $C_{t},\widetilde{C}_{t},i_{t},o_{t}$ and the weight matrices. For a detailed description, see [62]. Refer to Figure 2 for illustration of LSTM module.
@@
This model is known as triple exponential smoothing, where ETS stands for Error, Trend, and Seasonality terms. This model generalizes the Holt-Winters model by incorporating multiplicative error and trend terms. The term ”Exponential Smoothing” derives from the fact that the model’s equations can be expanded to represent an average of past observations with exponentially decaying weights. We chose this model for its reliability and widespread use, offering an excellent balance between performance and computational efficiency. For a detailed description and justification of this choice, see [16]. For short series, specifically those with fewer than 50 observations, ETS and its special case, Holt’s model, may be the most effective, as this quantity of data is typically sufficient to train models with reasonable accuracy; further details can be found in [16, 23]. When 50 or more observations are available, more complex models may be necessary, which we address using LSTM. In our experiments, we selected Holt’s model, which is additive version of ETS without seasonality, as it yielded the best results and has been proven effective in practice. We excluded seasonality because the series in our experiments generally lack seasonal variation. Let us briefly describe Holt’s model:
@@
where $L_{t},T_{t}$ represent the Level and Trend terms, respectively, and $\alpha,\gamma$ are smoothing parameters that typically fall within the range $(0,1)$. The prediction $h$ time steps ahead is given by
@@
The smoothing parameters can be computed using standard methods, such as minimizing the RSS for historical data.
@@
We included Linear Regression because it is one of the most widely used and most established statistical models and provides decent results for simple and regular time series. Polynomial Regression is analogous to Linear Regression but fits the data with a quadratic function instead of a linear one. Similar to Linear Regression, this model is straightforward and yields good results for simple and regular time series, particularly for short prediction horizons.
@@
This is a model that computes the historical growth rate $g$ on the training series using the following expression:
$g_{t}=\frac{1}{N}\sum_{t=1}^{N}y_{t}-y_{t-1}$
where $N$ is the number of observations. Predictions are generated by multiplying the last observation by the average growth rate and repeating this for successive forecast dates:
$\widetilde{y}_{t+h}=y(t)g(t)^{h}$
This simplistic method is used to enrich the ensemble.
@@ 3.2 Ensembling
As mentioned in the literature review, numerous articles demonstrated that ensembles are often more accurate than any single model on its own. The process of pooling models for the ensemble involves computing accuracy scores for each model. These models are then combined into a single prediction using a weighted arithmetic mean, as described by the formula:
@@
where $w_{i}$ represents the weights and $\widehat{y}_{i,t+h}$ denotes the forecasts at time step $t$ and $h$ steps ahead for the $i$-th model.
@@
The weights of the models are computed based on their accuracy scores, ensuring that more accurate models receive a higher weight. The metric employed is the Mean Absolute Error (MAE). We also experimented with other metrics, including Mean Squared Error, Root Mean Squared Error, and Mean Absolute Percentage Error. However, the MAE metric yielded the best performance in the final results.
Since higher metric values indicate greater errors, we invert these values so that a higher metric value reduces the model’s weight.
@@
$mae^{\prime}_{i}$ $=1-\frac{mae_{i}}{\sum_{i}mae_{i}},$
where $mae_{i},mae^{\prime}_{i},w_{i}$ represent the values of the metric, inverted metric, and weight for the $i$-th model, respectively. To compute the metric value, we separate 20% of the available dataset into test data. The model is trained on the reduced dataset, and forecasts are generated for the same dates as the testing data. The errors of the predictions relative to the testing data are computed using the metric, as expressed in the following equation:
$mae_{i}=\frac{1}{H}\sum_{h=1}^{H}\big{|}\widehat{y}_{i,t+h}-y_{t+h}\big{|},$ (3)
@@
See Figure 5 for a diagram illustrating the ensemble technique.
@@
The success of the ensemble technique lies in the fact that models in the ensemble tends to complement each other. Forecasts of well-performing models are amplified by assigning them greater weights, thereby exerting a more significant influence on the combined forecast. However, even high-quality forecasts have their flaws. In certain data segments, other models, despite having lower overall performance, may demonstrate more desirable behaviour. Thus, by averaging, these models enhance accuracy with a relatively minor effect on the strengths of the well-performing models, as they receive lower weights. In summary, through the process of averaging, this method has the capability to amplify the strengths, and with minimal impact to overall improvement, also mitigate the weaknesses, of individual models.
@@ 3.3 Direct Multi-Step (DMS)
In a typical DMS procedure, a forecast is first computed one step ahead. The resulting forecast is then added to the training data, and the model is retrained on this extended dataset. This procedure is repeated for the number of steps in the forecasting horizon. In our methodology, to decrease the amount of computing resources used, we refrain from applying DMS at each step. This is because retraining are the most computationally expensive part of forecast. Specifically, we retrain the forecasting models at two key intervals: once after the first year/month to generate 5-year/month forecasts, and again after five years/month to produce 20 to 30-year/month forecasts. These horizons correspond to the commonly recognized short and medium-term economic and financial forecasting periods. For the remaining steps, we employ the standard IMS procedure, which simply shifts the forecast input by one step without retraining the model or using ensembling. This limited retraining schedule allows us to retain the benefits of DMS, including its robustness to non-stationary data and structural changes, while significantly reducing the computational burden typically associated with full DMS implementations. Our methodology can be decomposed into the following three-step algorithm, where, for illustrative purposes, we specify the frequency to be annual with retraining at the first and fifth years.
@@
    Step 1 - One year ahead forecast - one-year-ahead forecasts are computed for each model in the ensemble. The ensemble technique described in subsection 3.2 is then applied. Specifically, the available data is divided into training and testing sets in an 80/20 ratio, and forecasts are computed for the testing observations. The MAE metric is used to compute performance weights according to formulas (2) and (3). Finally, the one-year forecasts are ensembled using the calculated weights, resulting in the ensembled forecast according to (1).
@@
    where $w_{i,1}$ are the computed weights and $\widetilde{y}_{i,t+1}$ are the forecasts at the time step one year ahead from time $t$ for the $i$-th model, which is trained on the training data up to time $t$, denoted as $Y_{t}=(y_{1},y_{2},\ldots,y_{N})$.
@@
    Step 2 - Retrain models and forecast five year ahead - the models are retrained with the training data extended by the ensembled forecast from Step 1. Five-year-ahead forecasts are then computed using the IMS procedure and ensembled in the same manner as in Step 1. The explicit formula for the forecast is given by the following expression:
    $\widehat{y}_{t+1+h}=\sum_{i}w_{5,i}\widetilde{y}_{t+5,i}(Y_{t},\widehat{y}_{t+1}),$
    where $h=1,2,3,4$ is the length of the forecasting horizon up to five years ahead, $w_{i,h}$ are the computed weights, and $\widetilde{y}_{t+h,i}$ are the forecasts at the time step five years ahead from time $t$, for the $i$-th model, trained on the initial training data $Y_{t}$ and the forecast from previous step.
@@
    Step 3 - Retrain and forecast full horizon - the model is retrained on the training data further extended by the ensembled forecasts from Step 2. Ensembled forecasts for the full horizon are then computed as previously described. The expression for the forecasts is as follows:
    $\widehat{y}_{t+5+h}=\sum_{i}w_{5+h,i}\widetilde{y}_{t+5+h,i}(Y_{t},\widehat{Y}_{t+5}),$
    where $h=1,2,\ldots,H-5$ is the length of the forecasting horizon up to horizon $H$, $w_{h,i}$ are the computed weights, and $\widetilde{y}_{t+5+h,i}$ are the forecasts for each model, trained on data further extended by forecasts up to the fifth year, denoted as $\widehat{Y}_{t+5}=(\widehat{y}_{t+1},\widehat{y}_{t+2},\ldots,\widehat{y}_{t+5})$.
@@
Typical DMS, as mentioned in the literature review, tends to improve performance, especially in medium- and long-term horizons. In contrast, IMS, even when combined with ensemble forecasting, tends to accumulate errors. Typical Decision-Making Systems (DMS), as highlighted in the literature review, generally enhance performance, particularly over medium- and long-term horizons. In contrast IMS, even when combined with ensemble forecasting, tend to accumulate errors. When initial forecasts contain relatively small errors, feeding these forecasts back into the same model to generate subsequent predictions can amplify those initial inaccuracies, as the model retains its inherent flaws. Given the substantial body of research on DMS, it is evident that these systems possess the capability to mitigate this accumulation of errors, with the reduction effect becoming increasingly significant as the forecasting horizon extends. This effect is especially visible in macroeconomic forecasting, encompassing time series such as GDP, inflation, and employment. Macroeconomic time series often exhibit nonlinear and non-stationary behaviour and are subject to structural breaks resulting from policy changes, economic shocks, or other external influences. DMS forecasting is better suited to handle these complexities because they do not depend on recursively generated inputs that may carry forward outdated or biased information. Instead, the procedure is designed to directly incorporate recent trends and shifts in the data.

### v2

@@ 3 Ensembled Direct Multi-Step forecasting
In this section, we describe in detail Ensembled Direct Multi-Step (EDMS) forecasting method. Subsection 2.1 presents models used for the ensemble forecasts, Subsection 2.2 discusses Ensembling method itself, and finally, Subsection 2.3 explores application of Direct Multi-Step (DMS) forecasting combined with Ensembling method.
@@ 3.1 Models
To enhance the modelling capability of the ensemble, we carefully selected a diverse set of models. We considered only models that perform well in terms of accuracy and computational efficiency. Additionally, we included models across full spectrum of complexity, ranging from simple to highly complex. For the simplest models, we employed basic forecasting using the average of past growth rates, as well as Linear and Polynomial Regression. For models of medium complexity, we utilized classical time series models, specifically General Exponential Smoothing (ETS). For the most complex model, we implemented an deep learning model, namely LSTM. In selecting these models, we aimed to minimize the computing resources and time required for forecasting, as economic and financial datasets tend to be relatively large. At the same time, we strive to maintain the highest possible accuracy. Prophet was excluded from the set of models due to significant computation time and gave minimal, if any, improvement in accuracy. We also have a constraint in the model selection that the datasets must be sufficiently large, as a considerable amount of data is necessary to train more complex models such as LSTM.
@@
    Performance forecast - these forecasts are calculated to assess the performance of the methods employed for computing weights in the ensembling process. We allocate 80% of the available data for training and 20% for testing. This split ratio was selected to ensure that there is sufficient data for effective model training, while also maintaining a reasonably large test set for proper model comparison. Experiments have demonstrated that the 80/20 ratio yields favourable results. Once forecasts are computed using the model trained on the training data, the value of performance metric is computed based on testing data
@@
    Full forecast - this forecasts are used to compute the ensembled forecast. We utilize 100% of the data for training and compute desired forecasts with the model trained on the complete dataset in order to get the best possible forecast.
Let us now individually describe each model that is used in the ensemble:
@@
This is a type of Recurrent Neural Network (RNN) model, which is a natural choice for modelling time series due to its ability to effectively recognize and learn patterns in sequential data. Notably, LSTM is well-known and widely used model for this purpose, as outlined in [62]. We chose this model for several reasons. First, according to a literature review in [50], LSTM is the most extensively studied deep learning model for time series forecasting, and there is a lot of information to support it’s usage. Particular reason for this selection is performance. Multiple sources indicate that LSTM outperforms widely used classical model ARIMA, as noted in [58, 44, 31, 72]. Furthermore, LSTM often prove to be the best performers among deep learning models, particularly for economic and financial data, as seen in [72, 59].
  The Long Short-Term Memory ANN architecture consists of sequentially connected modules with the number $n$ of these modules equal to the size of the input sequence $(y_{1},y_{2},\ldots,y_{n})$. In this sequence, each module takes the following inputs:
@@
  • Previous cell state $C_{t-1}$ - the cell state of the previous module, which facilitates long term memory.
@@
A single module is composed of the following components:
@@
    where $U_{f},W_{f}$ are weight matrices with weights parameters which determined in the training. This gate is depicted as first from left yellow rectangular in the module in Figure 2.
@@
    Update gate - this gate first computes a fully connected layer with hyperbolic tangent activation tanhtanh that proposes information $\widetilde{C}_{t}$ for the new cell state, where
    $\tanh(y)=\frac{\exp(y)-\exp(-y)}{\exp(y)+\exp(-y)}.$
    Update gate also computes a fully connected layer with sigmoid activation $\sigma$, known as the input layer $i_{t}$, selects which of the proposed information to add to the cell state. Both layers use previous output $\widetilde{y}_{t}$ and the element of the input sequence $y_{t}$ as inputs. Then this gate combines these two outputs to compute an update to the state $C_{t}$. Those computations can be expressed as follows:
@@
    where $U_{i},W_{i}$ and $U_{C},W_{C}$ are weight matrices, with weights parameters which determined in the training. This gate is depicted as two connected middle yellow rectangulars together with connection with Input Gate in the module in Figure 2.
@@
    Output gate - this final gate consists of a fully connected layer with sigmoid activation $\sigma$, with $\widetilde{y}_{t}$ and $y_{t}$ as inputs, which selects which parts of the cell state $C_{t}$, processed through $\tanh$ layer, to output :
@@
    where $U_{o},W_{o}$ are weight matrices with weights parameters which determined in the training. This gate is depicted as first from the right yellow rectangular in the module, together with connection to the output, in Figure 2.
The size and dimensions of $y_{t},\widetilde{y}_{t}$ are determined by the specific case at hand, which also determines the dimensions of $C_{t},\widetilde{C}_{t},i_{t},o_{t}$ and the size of weight matrices. For a detailed description, see [62]. Refer to Figure 2 for illustration of LSTM module.
@@
This model is known as triple exponential smoothing, where ETS stands for Error, Trend, and Seasonality terms. This is generalization the Holt-Winters model by incorporating multiplicative error and trend terms. The term ”Exponential Smoothing” derives from the fact that the model’s equations can be expanded to represent an average of past observations with exponentially decaying weights. We chose this model for its reliability and widespread use, offering an excellent balance between performance and computational efficiency. For a detailed description and justification of this choice, see [16]. For short series, specifically those with fewer than 50 observations, ETS and its special case, Holt’s model, may be the most effective, as this quantity of data is typically sufficient to train models with reasonable accuracy, further details can be found in [16, 23]. When 50 or more observations are available, more complex models may be necessary, we address that by using LSTM. In our experiments, we selected Holt’s model, which is additive version of ETS without seasonality, as it yielded the best results and has been proven very effective in term of computation time. We excluded seasonality because the series in our experiments generally lack seasonal variation. Let us briefly describe Holt’s model:
@@
where $L_{t},T_{t}$ represent the Level and Trend terms, respectively, and $\alpha,\gamma$ are smoothing parameters that typically fall within the range $(0,1)$. The prediction $h$ time steps ahead is given by:
@@
The smoothing parameters can be computed using standard methods, such as minimizing RSS for historical data. The starting values $L_{1}$ and $T_{1}$ are computed using first few observations in the series, e.g. $L_{1}=\sum_{i=1}^{s}x_{i}/2$.
@@
Linear Regression is one of the most widely used and most established statistical models. So we considered this model as a candidate for a model of low complexity. As expected, model yielded a decent results for simple and regular time series. Polynomial Regression is analogous to Linear Regression but fits the data with a quadratic function instead of a linear one. Similar to Linear Regression, this model is straightforward and yields good results for rather regular time series, and was particularly good for short prediction horizons.
@@
A simplistic method is used to enrich the ensemble. The method computes the historical growth rate on the training series using taking median of the relative changes of values from one observation to other, that is, the following expression:
$\delta(t)=1+\frac{y_{t}-y_{t-1}}{y_{t-1}}$
where $t$ is time of single observation. So the historical growth rate is a median of values $\widehat{\delta}=\text{median}(\delta(t),t=1,\ldots,N)$. Predictions are generated by multiplying the last observation by the average growth rate and repeating this for successive forecast dates:
$\widetilde{y}_{t+h}=y_{t}\widehat{\delta}^{h}$
where $h$ is the number of observation in the forecast horizon.
@@ 3.2 Ensembling
As mentioned in the literature review, numerous articles demonstrates that ensembles are often more accurate than any single model on its own. The process of pooling models for the ensemble involves computing accuracy scores for each model. These models are then combined into a single prediction using a weighted arithmetic mean, as described by the formula:
@@
where $w_{i}$ represents the weights and $\widehat{y}_{t+h,i}$ denotes the forecasts at time step $t$ and $h$ steps ahead for the $i$-th model.
@@
Since the weights of the models are computed based on accuracy scores, we can ensure more accurate models receive a higher weight. The metric employed in computation of the score is the Mean Absolute Error (MAE). We also experimented with other metrics, including Mean Squared Error, Root Mean Squared Error, and Mean Absolute Percentage Error. However, the MAE metric yielded the best performance in the final results.
Since higher metric values indicate greater errors, we need to invert these values so that a higher metric value reduces the model’s weight:
$mae^{\prime}_{i}=1-\frac{mae_{i}}{\sum_{i}mae_{i}},$
and so we have the following expressions for the weights:
@@
where $mae_{i},mae^{\prime}_{i},w_{i}$ represent the values of the metric, inverted metric, and weight for the $i$-th model, respectively. To compute the metric value, we separate 20% of the available dataset into test data. The model is trained on the dataset reduced to 80%, and forecasts are generated for the same dates as the testing data. The errors of the predictions relative to the testing data are computed using the metric, as expressed in the following expression:
$mae_{i}=\frac{1}{H}\sum_{h=1}^{H}\big|\widehat{y}_{t+h,i}-y_{t+h}\big|,$ (3)
@@
See Figure 5 for a diagram illustrating how the ensemble technique is carried out.
@@
The success of the ensemble technique lies in the fact that models in the ensemble tends to complement each other. Forecasts of well-performing models are amplified by assigning them greater weights, thereby exerting a more significant influence on the combined forecast. However, even high-quality forecasts have their flaws. In certain data segments, other models, despite having lower overall performance, may demonstrate more desirable behaviour. Thus, by averaging, these less-performing models enhance accuracy with a relatively minor effect on the strengths of the well-performing models, as they receive lower weights. In summary, through the process of averaging, this method has the capability to amplify the strengths, and with minimal impact to overall improvement, also mitigate the weaknesses, of individual models.
@@ 3.3 Direct Multi-Step (DMS)
In a typical DMS procedure, a forecast is first computed one step ahead. The resulting forecast is then added to the training data, and the model is retrained on this extended dataset. This procedure is repeated for the number of observations in the forecasting horizon. Retraining is the most computationally expensive part of forecast, hence, to decrease the amount of computing resources used, we refrain from applying DMS at each step. Specifically, we retrain the models at two key intervals for economic and financial data respectively: once after the first year/month to generate 5-year/month forecasts, and again after five years/month to produce 20 to 30-year/month forecasts. These horizons are can be seen as commonly recognized short and medium-term economic and financial forecasting periods. For the remaining steps, we employ the standard IMS procedure, which simply shifts the forecast input by one observation without retraining the model or using ensembling. This limited retraining schedule allows us to retain the benefits of DMS, including its robustness to non-stationary data and structural changes, while significantly reducing the computational burden typically associated with full DMS implementations. This methodology can be decomposed into the following three-step algorithm, where, for illustrative purposes, we specify the frequency to be annual with retraining at the first and fifth years:
@@
    Step 1 - One year ahead forecast - these forecast are computed for each model in the ensemble. The ensemble technique described in subsection 3.2 is then applied. Specifically, the available data is divided into training and testing sets in an 80/20 ratio, and forecasts are computed for the testing observations. The MAE metric is used to compute performance weights according to formulas (2) and (3). Finally, the one-year forecasts are ensembled using the calculated weights, resulting in the ensembled forecast according to (1).
@@
    where $w_{i,1}$ are computed weights and $\widetilde{y}_{t+1,i}$ are the forecasts at the time step one year ahead from time $t$ for the $i$-th model, which is trained on the training data up to time $t$, denoted as $Y_{t}=(y_{1},y_{2},\ldots,y_{t})$.
@@
    Step 2 - Retrain models and forecast five year ahead - the models are retrained with the training data extended by the ensembled forecast from Step 1. Then forecasts up to five-year-ahead horizon are then computed using the IMS procedure and ensembled in the same manner as in Step 1. The explicit formula for the forecast is given by the following expression:
    $\widehat{y}_{t+1+h}=\sum_{i}w_{5,i}\widetilde{y}_{t+1+h,i}(Y_{t},\widehat{y}_{t+1}),$
    where $h=1,2,3,4$ are the next time steps in the forecasting horizon, $w_{i,5}$ are computed weights for five year forecasts, and $\widetilde{y}_{t+1+h,i}$ are the forecasts at the time steps up five years ahead from time $t+1$, for the $i$-th model, trained on the initial training data $Y_{t}$ and the forecast from previous step.
@@
    Step 3 - Retrain and forecast full horizon - the model is retrained on the training data further extended by the ensembled forecasts from Step 2. Ensembled forecasts up to the full horizon are then computed as previously described. The expression for the forecasts is as follows:
    $\widehat{y}_{t+5+h}=\sum_{i}w_{H,i}\widetilde{y}_{t+5+h,i}(Y_{t},\widehat{Y}_{t+5}),$
    where $h=1,2,\ldots,H-5$ are the next time steps in the forecasting horizon, $w_{H,i}$ are the computed weights for full horizon, and $\widetilde{y}_{t+5+h,i}$ are the forecasts for each model, trained on data further extended by forecasts up to the fifth year, denoted as $\widehat{Y}_{t+5}=(\widehat{y}_{t+1},\widehat{y}_{t+2},\ldots,\widehat{y}_{t+5})$.
@@
Typical DMS, as highlighted in the literature review, generally enhance performance, particularly over medium- and long-term horizons. In contrast IMS, even when combined with ensemble forecasting, tend to accumulate errors. When initial forecasts contain relatively small errors, feeding these forecasts back into the same model to generate subsequent predictions can amplify those initial inaccuracies, as the model retains its inherent flaws. Given the substantial body of research on DMS, it is evident that this procedure possess the capability to mitigate this accumulation of errors, with the reduction effect becoming increasingly significant as the forecasting horizon extends. This effect is especially visible in macroeconomic forecasting, encompassing time series such as GDP, inflation, and employment over long horizons. Macroeconomic time series often exhibit non-linear and non-stationary behaviour and are subject to structural breaks resulting from policy changes, economic shocks, or other external influences. DMS forecasting is better suited to handle these complexities because it do not depend on recursively generated inputs that may carry forward outdated or biased information. Instead, the procedure is designed to directly incorporate recent trends and shifts in the data.